OpenStudy (anonymous):

Evaluate the velocity of a body and its distance moved, both after 2 minutes from it starts from rest, when its acceleration is given by a = 15 m/s^2

OpenStudy (anonymous):

Clicked closed by accident on last post :/

OpenStudy (anonymous):

Any help? :( this if for calculus study

OpenStudy (welshfella):

acceleration = dv/dt = 15 constant separating variables dv = 15dt integrating v = 15t + c when t = 0 v = 0 ( as it starts from rest) so c = 0 v = 15 t so substitute t = 120 to find the velocity after 2 minutes

OpenStudy (welshfella):

ok?

OpenStudy (anonymous):

So v = 120 x 15 =1800 m/s?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Now how will I go about finding the distance moved?

OpenStudy (anonymous):

Ps thanks for the help - i struggle with calculus haha

OpenStudy (welshfella):

a further integration will give you the distance travelled in terms of t the constant of integration will again be 0 because it was a start from rest

OpenStudy (welshfella):

cwrw238 asked me to look at this for you as she had to go to sit an examination

OpenStudy (welshfella):

integrating will give you distance = 15 * t^2/2

OpenStudy (anonymous):

So is the answer 108000m^2 or 24300000m^2 :/

OpenStudy (anonymous):

Oh that was nice of her and nice of you to help :)

OpenStudy (welshfella):

she made a mistake she put t = 1800 instead of t = 120 answer is 1/2 * 15 * 120^2 = 108,000 meters

OpenStudy (welshfella):

i'll kid her about that when i see her next!! its not like her to make mistakes - but we all do!

OpenStudy (anonymous):

Ahh I see that's what was getting me soo confused lol. That's true! Haha well thank you both for your help!! I'm so happy to have got that question done :) :)

OpenStudy (welshfella):

yw